Please help me with this question and show the steps of how you got it:
Jill buys 10 shares of stock in Costco Corporation at $150 per share on January 30, a membership retailer. During the year, Costco had a dividend of $1.80 per share at the end of June and the stock rose to $164.00 a share. If Jill sold her stock on July 31, what would be her return?
1
Expert's answer
2020-04-13T16:48:42-0400

decide by the formula of current profitability

"r=\frac{d}{P}"


find stock returns

"r=\frac{d}{P}=\frac{\$1.8}{150}=\$0.012,"


find dividends

"d=r\times P=\$164*0.012=\$1.968."